A car starts from rest and accelerates uniformly for a five seconds along a straight road. If speed obtained by the car is 72 km/h, then its acceleration is 


1
Expert's answer
2021-08-04T16:24:30-0400

By definition, the acceleration is:


"a = \dfrac{v_f - v_i}{t}"

where "v_f = 72km\/h = 20m\/s" is the final speed, and "v_i = 0" is the initial speed (start from rest). "t = 5s" is the time.

Thus, obtain:


"a = \dfrac{20m/s - 0m/s}{5s} = 4m/s^2"

Answer. 4 m/s^2.